# TESGNN: Temporal Equivariant Scene Graph Neural Networks for Efficient and Robust Multi-View 3D Scene Understanding

### v2

@@ IV-A Graph Initialization
The node feature includes the invariant features $\mathbf{h_{i}}$ and the 3-vector coordinate $\mathbf{x_{i}}\in\mathbb{R}^{3}$. The invariant features $\mathbf{h}_{i}$ consists of the latent feature of the point cloud after going through the PointNet $f_{p}(\mathbf{P_{i}})$, standard deviation $\boldsymbol{\sigma}_{i}$, log of the bounding box size $\ln\left(\mathbf{b}_{i}\right)$ where $\mathbf{b}_{i}=\left(b_{x},b_{y},b_{z}\right)\in\mathbb{R}^{3}$, log of the bounding box volume $\ln(v_{i})$ where $v_{i}=b_{x}b_{y}b_{z}$, and log of the maximum length of bounding box $\ln(l_{i})$. The coordinate of the bounding box $x_{i}$ is defined by the coordinates of the two furthest corners of the bound box. The formula can be written as follows:
@@
$\mathbf{h}_{i}$ $=\left[f_{p}\left(\mathbf{P}_{i}\right),\boldsymbol{\sigma}_{i},\ln\left(\mathbf{b}_{i}\right),\ln\left(\nu_{i}\right),\ln\left(l_{i}\right)\right],$
@@
$\begin{gathered}\mathbf{r}_{ij}=\left[\overline{\mathbf{p}}_{i}-\overline{\mathbf{p}}_{j},\boldsymbol{\sigma}_{i}-\boldsymbol{\sigma}_{j},\mathbf{b}_{i}-\mathbf{b}_{j},\ln\left(\frac{l_{i}}{l_{j}}\right),\ln\left(\frac{\nu_{i}}{\nu_{j}}\right)\right],\\ \mathbf{e}_{ij}=g_{s}\left(\mathbf{r}_{ij}\right),\end{gathered}$
@@
IV-B Graph Neural Network Architecture and Weights Updating
Fig. 3 describes our GNN network with two core components: ① Feature-wise Attention Graph Convolution Layer (FAN-GCL); and ② Equivariant Graph Convolution Layer (EGCL). The FAN-GCL, proposed by [11], is used to handle the large input queries $Q$ of dimensions $d_{q}$ and targets $T$ of dimensions $d_{\tau}$ by utilizing multi-head attention. On the other hand, EGCL, proposed by [19], is used to maintain symmetry-preserving equivariance, allowing us to incorporate the bounding box coordinates $x_{i}$ as node features and update them through the message-passing mechanism.
ESGNN has 2 main layers, each consists of a FAN-GCL followed by an EGCL, forming a total of 4 message-passing layers. For each main layer, the formulas for updating node and edge features are defined as:
@@ IV-B Graph Neural Network Architecture and Weights Updating
V Temporal Graph Matching Network
Figure 4: Temporal Graph Matching pipeline. First, node and edge embeddings derived from scene graphs of different sequences (a). For each sequence, edge embeddings are concatenated with the target node embeddings to create Predicate-Object embeddings (b), which then pass through a linear layer followed by sum pooling (c). For each node, the embeddings from all associated edges are concatenated and processed through another linear layer and self-attention mechanism to generate the final representation of each segment (d). These final embeddings are utilized for top-K retrieval graph matching (e).
By leveraging the symmetry-preserving properties of our equivariant network, we hypothesize that its node and edge embeddings can remain inherently distinguishable, especially for isomorphic scene graphs subjected to rotations and translations. This unique feature opens the gate for an interesting application: Given sequences of point clouds capturing overlapping regions of the same scene, our model can reliably match the corresponding parts, regardless of viewpoint changes. As a result, it facilitates the merging of multiple scene graphs into a unified global representation, ensuring spatial consistency and completeness over time.
Figure 4 illustrates our graph matching. Since sequences may differ in the number of nodes and edges, we define a Node-Edge representation for each node and its connected edges. Our goal is to produce consistent representations for identical node-edge embeddings appearing across sequences.
[…]
Contrastive Loss Our training goal is to optimize the embedding model such that distances between similar node-edge embeddings are minimized, while those between dissimilar embeddings are maximized. We achieve this using Contrastive Loss [37], a well-established method for training embeddings [38]. Unlike traditional loss functions that aggregate over samples, Contrastive Loss operates on pairs:
@@ V Temporal Graph Matching Network
where $(Y,X_{1},X_{2})$ are pairs of node-edge embeddings, $D_{W}$ is a distance function, and $m>0$ is a margin. We use Siamese Cosine Similarity [38] for $D_{W}$ and train with Online Contrastive Loss, focusing on hard positive and hard negative pairs. These are widely-used methods in Sentence Transformer [38], which inspire our approach.
@@
Figure 5: Comparison of recall result over training steps, column-wise interpretation. (a), (d) illustrates the result of our model versus SGFN. (b), (e) shows the ablation study. (c), (f) shows result while applying our model with the image encoder.
VI Experiment
We evaluate TESGNN on the 3DSSG dataset [10] and compare the results with state-of-the-art works. Section VI-C provides result for scene graph generation from 3D point clouds, in comparison to 3DSSG [10] and Scene Graph Fusion (SGFN) [11]. Our method is evaluated on full scenes given geometric segments mentioned in Section III-B. Section VI-E reports our Temporal Graph Matching.
VI-A Dataset
Scene Graph Extraction: We use the 3DSSG [10]  [^1] https://github.com/ShunChengWu/3DSSG. The scene graph ground truth in 3DSSG contains 2 versions: l20, which includes 20 objects and 8 predicates, and l160, which includes 160 objects and 26 predicates. We mainly use the test set of the l20 version for our experiment as the l160 is currently having some issues within the dataset, as confirmed by the authors. - a popular dataset for 3D scene graph evaluation built upon 3RScan[35], adapting the setting from SGFN [11]. The original 3RScan contains 1482 3D reconstructions/snapshots of 478 naturally changing indoor environments. After processed with ScanNet [39] for geometric segmentation and ground truth scene graph generation, the final dataset consists of 1061 sequences from 382 scenes for training, 157 sequences from 47 scenes for validation, and 117 sequences from 102 scenes for test.
Temporal Graph Matching: While reproducing prior works including SGAligner [29] and SG-PGM [30], we identified limitations in their dataset. Since their work does not contain scene graph extraction, they synthesize scene graph node and edge features using one-hot encoding of the ground truth labels, which does not accurately reflect scene graph embeddings. This set-up impractically ignores real-world noises affecting local scene graphs. To address these shortcomings, we leverage our above set-up for scene graph extraction to evaluate scene graph matching, using the same sequences and scenes. In this context, a positive pair consists of the same object appearing in two or more different sequences. There are 16,156 positive pairs over 321,292 total pairs for train set, and 3,062 positive pairs over 63,890 total pairs for test set. In our set-up, node and edge features are derived from frozen scene graphs such as ESGNN (ours) or SGFN [11].
VI-B Metrics
[…]
VI-C Scene Graph Extraction Evaluation
Better prediction outcome: Table I compares the results between ESGNN (ours) with existing models 3DSSG [10] and SGFN [11] on the 3DSSG-l20 dataset with geometric segmentation setting. Our method obtains high results in both relationship, object, and predicate classification. Especially, ESGNN outperforms the existing methods in relationship prediction and obtains significantly higher R@1 in predicate compared to SGFN. For unseen data, ESGNN is competitive with SGFN, as shown in Table II.
@@ VI-C Scene Graph Extraction Evaluation
3DSSG 32.65 50.56 55.74 83.89 $\mathbf{95.22}$ 98.29 55.74 $\mathbf{95.22}$
SGFN 37.82 48.74 62.82 $\mathbf{88.08}$ 81.41 98.22 63.98 94.24
Ours $\mathbf{43.54}$ $\mathbf{53.64}$ $\mathbf{63.94}$ 86.65 94.62 $\mathbf{98.30}$ $\mathbf{65.45}$ 94.62
TABLE I: Scene graph predictions for relationship triplet, object, and predicate, measured on 3DSSG-l20. The Recall column reports the recall scores on objects (Obj.) and relationships (Rel.)
@@
3DSSG 39.74 49.79 55.89 84.42 $\mathbf{70.87}$ 83.29
SGFN $\mathbf{47.01}$ 55.30 64.50 $\mathbf{88.92}$ 68.71 $\mathbf{83.76}$
Ours 46.85 $\mathbf{56.95}$ $\mathbf{65.47}$ 87.52 66.90 82.88
TABLE II: Scene graph predictions for new unseen relationship triplet, object, and predicate, measured on 3DSSG-l20 with geometric segmentation.
@@
More robust against imbalanced data: We additionally provide the training confusion matrices after the first 1000 steps in Figure 6. It is notably observable that ESGNN 6(a), 6(c) shows clearer diagonal patterns compared to SGFN 6(b), 6(d), which was biased towards some dominant classes. These patterns demonstrate that ESGNN is more data-efficient than SGFN, as it does not need to generalize over rotations and translations of the data, while still harnessing the flexibility of GNNs in larger datasets. However, due to its high bias [19], our EGCL-based message passing performs well with limited data but struggles to learn the dataset’s subtleties as the training size increases. In later training steps, we observe the overfitting. Yet, our model consistently outperforms the pioneering works overall.
Scene Graph Extractor with Image Encoder: ESGNN are also potential to apply with point-image encoder instead of 3D point clouds only. Figures 5(c) and 5(f) show the performance of our implementation with image encoders, called Joint-ESGNN, compared to existing methods including JointSSG [12] and SGFN.
VI-D Ablation Study
Table III evaluates of ESGNN with different architectures and settings that we experimented. ⓪ is the SGFN, run as the baseline model for comparison. ① is the ESGNN with 1 FAN-GCL layer and 1 EGCL layer. This is our best performer and is used for experiments in Section VI-C. ② is similar to ①. The only difference is that we concatenate coordinate embedding to the output edge embedding after message passing. ③ and ④ are similar to ① and ② respectively, but with 2 FAN-GCL layers and 2 EGCL layers.

### v3

@@ 4.1 Graph Initialization
The node feature includes the invariant features $\mathbf{h_{i}}$ and the 3-vector coordinate $\mathbf{x_{i}}\in\mathbb{R}^{3}$. The invariant features $\mathbf{h}_{i}$ consists of the latent feature of the point cloud after going through the PointNet $f_{p}(\mathbf{P_{i}})$, standard deviation $\bm{\sigma}_{i}$, log of the bounding box size $\ln\left(\mathbf{b}_{i}\right)$ where $\mathbf{b}_{i}=\left(b_{x},b_{y},b_{z}\right)\in\mathbb{R}^{3}$, log of the bounding box volume $\ln(v_{i})$ where $v_{i}=b_{x}b_{y}b_{z}$, and log of the maximum length of bounding box $\ln(l_{i})$. The coordinate of the bounding box $x_{i}$ is defined by the coordinates of the two furthest corners of the bound box. The formula can be written as follows:
@@
$\mathbf{h}_{i}$ $=\left[f_{p}\left(\mathbf{P}_{i}\right),\bm{\sigma}_{i},\ln\left(\mathbf{b}_{i}\right),\ln\left(\nu_{i}\right),\ln\left(l_{i}\right)\right],$
@@
$\begin{gathered}\mathbf{r}_{ij}=\left[\overline{\mathbf{p}}_{i}-\overline{\mathbf{p}}_{j},\bm{\sigma}_{i}-\bm{\sigma}_{j},\mathbf{b}_{i}-\mathbf{b}_{j},\ln\left(\frac{l_{i}}{l_{j}}\right),\ln\left(\frac{\nu_{i}}{\nu_{j}}\right)\right],\\ \mathbf{e}_{ij}=g_{s}\left(\mathbf{r}_{ij}\right),\end{gathered}$
@@
4.2 Graph Neural Network Architecture and Weights Updating
Fig. 3 describes our GNN network with two core components: ① Feature-wise Attention Graph Convolution Layer (FAN-GCL); and ② Equivariant Graph Convolution Layer (EGCL). The FAN-GCL, proposed by Wu et al. (2021b), is used to handle the large input queries $Q$ of dimensions $d_{q}$ and targets $T$ of dimensions $d_{\tau}$ by utilizing multi-head attention. On the other hand, EGCL, proposed by Satorras et al. (2021), is used to maintain symmetry-preserving equivariance, allowing us to incorporate the bounding box coordinates $x_{i}$ as node features and update them through the message-passing mechanism.
ESGNN has 2 main layers, each consisting of a FAN-GCL followed by an EGCL, forming a total of 4 message-passing layers. For each main layer, the formulas for updating node and edge features are defined as:
@@ 4.2 Graph Neural Network Architecture and Weights Updating
Equivariance of ESGNN
We provide a detailed proof of equivariance in the Appendix B.
5 Temporal Graph Matching Network
Leveraging the symmetry-preserving properties of our scene graph extractor, we hypothesize that its node and edge embeddings can remain inherently distinguishable, especially for isomorphic scene graphs subjected to rotations and translations. This unique feature opens the gate for efficient temporal scene graph matching: Given multiple point cloud sequences containing overlapping regions, our model can reliably match the corresponding parts, regardless of viewpoint changes. As a result, multiple scene graphs can be aligned and fused into a unified, spatially consistent global graph over time.
Figure 4 illustrates our graph matching model. The core idea is to generate a compact representation for each unique triplet $\left<\textit{subject, predicate, object}\right$>, then merge identical ones by their similarity. This approach avoids explicit graph isomorphism search, while being designed to ensure order-invariant and scale to graphs of any size.
Figure 4: Temporal Graph Matching pipeline. First, node and edge embeddings derived from scene graphs of different sequences (a). For each sequence, edge embeddings are concatenated with the target node embeddings to create Predicate-Object embeddings (b), which then pass through a linear layer followed by sum pooling (c). For each node, the embeddings from all associated edges are concatenated and processed through another linear layer and self-attention mechanism to generate the final representation of each segment (d). These final Triplet Embeddings are utilized for top-K retrieval graph matching (e).
Triplet Representation For each triplet $\left<\textit{subject, predicate, object}\right$> we:
  • Concatenate the predicate (edge) embedding with the Object-node embedding to produce a Predicate–Object vector.
  • Pass the Predicate–Object vectors connected to the same Subject through a linear layer and sum-pooled to ensure permutation invariance.
  • Fuse the pooled vector with the Subject-node embedding, then refine it through a second linear layer and a self-attention block to capture higher-order context.
The result is a compact, rotation-invariant Triplet embedding for every local scene graph.
Similarity-based Graph Matching Given two sequences, we compute cosine similarities between all Triplet embeddings and perform top-K retrieval. Pairs whose similarity exceeds a fixed threshold are treated as the same physical object; the corresponding local graphs are merged node by node to build the global scene graph.
Training Objective We train the graph matching model such that distances between similar Triplet embeddings are minimized, while those between dissimilar embeddings are maximized. We achieve this using Contrastive Loss Hadsell et al. (2006), a well-established method for training embeddings Reimers & Gurevych (2019). Unlike traditional loss functions that aggregate over samples, Contrastive Loss operates on pairs:
@@ 5 Temporal Graph Matching Network
where $(Y,X_{1},X_{2})$ are pairs of Triplet Embeddings, $D_{W}$ is a distance function, and $m>0$ is a margin. We use Siamese Cosine Similarity Reimers & Gurevych (2019) for $D_{W}$ with hard-positive / hard-negative mining to focus learning on the most ambiguous cases.
6 Experiment
We evaluate TESGNN on the 3DSSG dataset Wald et al. (2020) and compare the results with state-of-the-art works. Section 6.3 provides results for scene graph generation from 3D point clouds, in comparison to 3DSSG Wald et al. (2020) and Scene Graph Fusion (SGFN) Wu et al. (2021b). Our method is evaluated on full scenes given geometric segments mentioned in Section 3.2. Section 6.5 reports our Temporal Graph Matching.
6.1 Dataset
Scene Graph Extraction: We use the 3DSSG Wald et al. (2020)  [^1] https://github.com/ShunChengWu/3DSSG - a popular dataset for 3D scene graph evaluation built upon 3RScan Wald et al. (2019), adapting the setting from SGFN Wu et al. (2021b). At the time this paper is written, 3DSSG is the only dataset for semantic scene graph generation. The original 3RScan contains 1482 3D reconstructions/snapshots of 478 naturally changing indoor environments. After being processed with ScanNet Dai et al. (2017) for geometric segmentation and ground truth scene graph generation, the final dataset consists of 1061 sequences from 382 scenes for training, 157 sequences from 47 scenes for validation, and 117 sequences from 102 scenes for testing. We tested TESGNN on both 2 dataset versions l20 and l160. The l20 contains 20 objects and 8 predicates, and l160 contains 160 objects and 26 predicates.
Temporal Graph Matching: While reproducing prior works including SGAligner Sarkar et al. (2023) and SG-PGM Xie et al. (2024), we identified limitations in their dataset. Since their work does not contain scene graph extraction, they synthesize scene graph node and edge features using one-hot encoding of the ground truth labels, which does not accurately reflect scene graph embeddings. This setup impractically ignores real-world noises affecting local scene graphs. To address these shortcomings, we leverage our above setup for scene graph extraction to evaluate scene graph matching, using the same sequences and scenes. In this context, a positive pair consists of the same object appearing in two or more different sequences. There are 16,156 positive pairs over 321,292 total pairs for the train set, and 3,062 positive pairs over 63,890 total pairs for the test set. In our setup, node and edge features are derived from frozen scene graphs such as ESGNN (ours) or SGFN Wu et al. (2021b).
@@ 6.1 Dataset
Figure 5: Comparison of recall result over training steps, column-wise interpretation. (a), (d) illustrates the result of our model versus SGFN. (b), (e) shows the ablation study. (c), (f) shows the result while applying our model with the image encoder.
6.2 Metrics
Scene Graph Extraction: We mainly use the Recall of node and edge as our evaluation metrics, given that the dataset is unbalanced Wald et al. (2020) and the objective of scene graphs is to effectively capture the semantic meaning of the surrounding world. In the training phase, we calculate the Recall as the true positive overall positive prediction. In the test phase, for more detailed analysis, we use Recall@k (R@k) Wald et al. (2020); Wu et al. (2021b; 2023), which takes k most confident predictions and marks them as correct if at least one of these predictions is correct. For a relationship triplet, R@1 is the accuracy of the predictions. We apply the recall metrics for the predicate (edge classification), object (node classification), and relationship (triplet $\left<\textit{subject, predicate, object}\right$>). Additionally, we report the Number of Params and Runtime for inference.
Temporal Graph Matching: We follow the Information Retrieval evaluation metrics, including R@k, MRR@k, and mAP@k  [^2] Since each node has only one ground-truth match, mAP@k and MRR@k produce similar result in our setting.. The R@k used in this evaluation is slightly different: It is the proportion of instances where at least one correct match appears among the topkk matches with a confidence score above 0.5. If the matching score is below 0.5, the node is considered unique, indicating that it does not have a common match across sequences. MRR@k evaluates the effectiveness of retrieval by averaging the reciprocal ranks of the $k$ similar nodes. mAP@k measures the quality of rankings by emphasizing $k$ relevant nodes at higher ranks.
6.3 Scene Graph Extraction Evaluation
@@
3DSSG 32.65 50.56 55.74 83.89 95.22 98.29 55.74 95.22
SGFN 37.82 48.74 62.82 88.08 81.41 98.22 63.98 94.24
Ours 43.54 53.64 63.94 86.65 94.62 98.30 65.45 94.62
Method Relationship Object Predicate Recall R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 Obj. Rel. 3DSSG 64.80 68.69 69.91 27.33 60.82 73.59 67.25 96.72 98.63 27.33 21.65 SGFN 64.69 68.22 69.37 35.76 67.55 79.40 48.63 92.33 97.71 35.76 14.09 Ours 65.00 69.38 70.46 37.46 67.97 79.14 36.88 90.14 96.89 37.46 20.91
Table 1: Scene graph prediction performance on l20 and l160 datasets, including relationship triplets, objects, and predicates. Recall columns reporting object and relationship recall scores (Obj., Rel.). As the l160 dataset contains more objects compared to l20, we provide a wider range of R@k for better assessment.
Effective prediction outcome: Tables 6(a) and 6(b) compare the results between ESGNN (ours) with existing models 3DSSG Wald et al. (2020) and SGFN Wu et al. (2021b) with a geometric segmentation setting. Our method obtains high results in both relationship, object, and predicate classification. Especially, ESGNN outperforms the existing methods in relationship prediction and obtains significantly higher R@k in predicate compared to SGFN. For the l160 dataset, while TESGNN maintains consistently stronger performance in relationship and object prediction, we observe a decrease in predicate prediction compared to existing methods. We attribute this to the influence of equivariance when distinguishing directional relationships (e.g., "left" vs. "right"), which appear more frequently in the l160 than in l20. Additionally, the authors of SGFN and 3DSSG suggested that the higher predicate recall in 3DSSG may come from the use of union 3D bounding boxes, which could be more suitable for estimating multiple predicates Wu et al. (2021a). Nevertheless, we highlight the high accuracy in relationship triplet prediction of our approach. This is crucial, as it demonstrates that ESGNN is effective in capturing semantic relationships between objects, a key factor in generating accurate scene graphs. For predictions on unseen data, ESGNN performs competitively with SGFN, as shown in Tables 7(a) and 7(b).
@@
3DSSG 39.74 49.79 55.89 84.42 70.87 83.29
SGFN 47.01 55.30 64.50 88.92 68.71 83.76
Ours 46.85 56.95 65.47 87.52 66.90 82.88
Method New Relationship New Object New Predicate
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
3DSSG 06.81 16.83 20.10 27.24 61.12 73.90 67.25 96.73 98.63
SGFN 06.70 15.59 18.65 35.51 67.38 79.35 48.62 92.32 97.70
Ours 07.84 18.69 21.55 37.22 67.78 79.05 36.87 90.13 96.87
Table 2: Scene graph prediction performance on previously unseen relationship triplets, objects, and predicates under geometric segmentation setting.
@@
Robustness against imbalanced data: We additionally provide the training confusion matrices after the first 1000 steps in Figure 8. It is notably observable that ESGNN 8(a), 8(c) shows clearer diagonal patterns compared to SGFN 8(b), 8(d), which was biased towards some dominant classes. These patterns demonstrate that ESGNN is more data-efficient than SGFN, as it does not need to generalize over rotations and translations of the data, while still harnessing the flexibility of GNNs in larger datasets. However, due to its high bias Satorras et al. (2021), our EGCL-based message passing performs well with limited data but struggles to learn the dataset’s subtleties as the training size increases. Additionally, in later training steps, we observe sign of overfitting. We anticipate that this behavior aligns with the EGCL design that intentionally allows overfitting to fully fit the training graphs Satorras et al. (2021), in which its authors also conducted a thorough overfitting evaluation. Section 6.4 further discusses our experiments to mitigate this effect. Nevertheless, our model consistently outperforms the existing methods overall.
@@
Computation Efficiency: Table 3 reports the model parameters and runtime. In this benchmark, we run the inference for each method 1000 times and calculate the average runtime. All experiments were conducted on an AMD Threadripper Pro 3955WX 16-core CPU and an NVIDIA RTX 4090 24GB GPU. The results show that ESGNN achieves better computational efficiency than existing methods. On CPU, ESGNN requires 26% fewer parameters and achieves a 1.1× speedup over SGFN, with higher FPS. For broader context, OpenFusion Yamazaki et al. (2024) reports 50 FPS for 3D semantic map reconstruction on the ScanNet dataset. Although their setup differs from ours and hardware specifications were not reported, we provide this information as a useful reference for understanding the runtime efficiency of scene graph generation and VLM-based semantic map reconstruction.
Method Params CPU GPU GNN Total Runtime FPS Runtime FPS 3DSSG 2,629,120 3,245,204 57.35 ms 17.44 (10× $\downarrow$) 2.51 ms 398.00 (1.2× $\downarrow$) SGFN 2,245,312 2,861,396 6.27 ms 159.48 (1.1× $\downarrow$) 2.20 ms 455.00 (1.1× $\downarrow$) Ours 1,780,577 2,396,661 5.61 ms 178.20 2.12 ms 471.78
Table 3: Performance comparison of different methods for scene graph generation, evaluated on both CPU and GPU. The reported runtime excludes the segmentation process and reflects only the scene graph generation phase, including node and edge prediction. $\downarrow$ denotes × times slower compared to ours.
Scene Graph Extractor with Image Encoder: ESGNN is also potentially applicable with a point-image encoder instead of 3D point clouds only. Figures 5(c) and 5(f) show the performance of our implementation with image encoders, called Joint-ESGNN, compared to existing methods, including JointSSG Wu et al. (2023) and SGFN.
6.4 Ablation Study
Table 4 evaluates ESGNN with different architectures and settings that we experimented with. ⓪ is the SGFN, run as the baseline model for comparison. ① is the ESGNN with 1 FAN-GCL layer and 1 EGCL layer. This is our best performer and is used for experiments in Section 6.3. ② is similar to ①. The only difference is that we concatenate the coordinate embedding to the output edge embedding after message passing. ③ and ④ are similar to ① and ②, respectively, but with 2 FAN-GCL layers and 2 EGCL layers.